# Sparse-to-Sparse Training of Diffusion Models

### v1

@@ 3.3.2 Experimental Details
To be able to compare the performance of different methods and different sparsity levels, we train the models for a predefined amount of epochs: 150 for Latent Diffusion datasets, and 600 for ChiroDiff datasets. For a complete description of training details please refer to Section B.3. Given the extensive number of experiments we conducted, we opted for a shorter training regime. For sampling, we use DDIM sampling [Song et al., 2021] with 100 steps for Latent Diffusion, and 50 steps for ChiroDiff, following the guidance provided in the original papers.
@@ 3.3.3 Evaluation Metrics
We follow common practice and calculate the FID score [Heusel et al., 2017] to assess the performance of all models. Refer to Section B.4 for more information on FID calculation. To evaluate the computational savings of the sparse methods, we report the network size (number of parameters) as a proxy for memory requirement, and the FLOPs, to estimate the computational cost of training and inference. We follow the method of FLOPs calculation described by Evci et al. [2020].
@@ 4 Experimental Results
We analyze the performance of Static-DM, MagRan-DM, and RigL-DM across various sparsity levels, and compare the results against the original Dense baseline. Later on, in Section 4.3 we present experiments comparing a selection of DST vs. Dense models across various diffusion timesteps. Examples of the generated samples can be found in Appendix G.
@@ 4.1 Latent Diffusion
Static-DM, $S=0.5$ 33.19 $\pm$ 2.39 0.50$\times$ 0.68$\times$ 0.68$\times$
RigL-DM, $S=0.25$ 32.12 $\pm$ 3.10 0.75$\times$ 0.91$\times$ 0.91$\times$
MagRan-DM, $S=0.5$ 32.83 $\pm$ 1.68 0.50$\times$ 0.67$\times$ 0.67$\times$
@@
Static-DM, $S=0.25$ 28.79 $\pm$ 12.65 0.75$\times$ 0.91$\times$ 0.91$\times$
RigL-DM, $S=0.10$ 37.80 $\pm$ 13.55 0.90$\times$ 0.97$\times$ 0.97$\times$
MagRan-DM, $S=0.25$ 28.20 $\pm$ 7.64 0.75$\times$ 0.91$\times$ 0.91$\times$
@@
Static-DM, $S=0.10$ 119.92 $\pm$ 5.94 0.90$\times$ 0.97$\times$ 0.97$\times$
RigL-DM, $S=0.10$ 121.59 $\pm$ 6.91 0.90$\times$ 0.97$\times$ 0.97$\times$
MagRan-DM, $S=0.25$ 117.32 $\pm$ 8.52 0.75$\times$ 0.91$\times$ 0.91$\times$
@@
In all datasets, Static-DM has better performance than the dynamic methods in higher sparsity setups, $S>0.5$. This is interesting, as it departs from the usual patterns found in sparse-to-sparse training for supervised learning applications and even other generative models such as GANs, where DST usually outperforms SST [Mocanu et al., 2018, Liu et al., 2023b]. Liu et al. [2021c] found that, in image classification tasks, DST models consistently achieve better performance over SST with appropriate parameter exploration, i.e., exploration frequency $\Delta T_{e}$ and prune and regrowth ratio $p$. To provide insights on the importance of $p$ for DST experiments, we conducted an experiment using a prune and regrowth rate $p\in\{0.05,0.1,0.2,0.3,0.5\}$. The results are provided in Figure 5 in Appendix F. The best results were obtained with $p=0.05$.
Following this experiment, we repeated all experiments for DST methods presented in Figure 2, using $p=0.05$, and show the results in Figure 6 and Appendix F. One particularly interesting finding is that, in high sparsity regimes, such as $S=0.9$ and $S=0.75$, DST methods have consistently better performance when $p=0.05$, even outperforming Static-DM. However, this performance advantage disappears when using the more aggressive prune and regrowth rate of $p=0.5$. Please refer to Appendix F for a more in-depth analysis.
@@ 4.2 ChiroDiff
Figure 3 shows the FID scores of the studied sparse methods for ChiroDiff. For QuickDraw, we observe that both Static-DM and RigL-DM exhibit variations around the performance of the Dense model, with only a subtle tendency to deteriorate as sparsity increases. MagRan-DM consistently matches the FID of the Dense model, and is able to outperform it at 90% sparsity. These results suggest that this model is overparameterized, which would explain why it benefits significantly from sparsity, even when removing 90% of the weights.
@@
Static-DM, $S=0.25$ 29.39 $\pm$ 0.24 0.75$\times$ 0.75$\times$ 0.75$\times$
RigL-DM, $S=0.10$ 29.38 $\pm$ 0.27 0.89$\times$ 0.89$\times$ 0.89$\times$
MagRan-DM, $S=0.90$ 29.45 $\pm$ 0.39 0.10$\times$ 0.10$\times$ 0.10$\times$
@@
Static-DM, $S=0.5$ 22.36 $\pm$ 0.87 0.50$\times$ 0.51$\times$ 0.51$\times$
RigL-DM, $S=0.25$ 21.14 $\pm$ 0.71 0.70$\times$ 0.70$\times$ 0.70$\times$
MagRan-DM, $S=0.25$ 21.73 $\pm$ 1.18 0.39$\times$ 0.39$\times$ 0.39$\times$
@@
Static-DM, $S=0.25$ 47.29 $\pm$ 1.96 0.75$\times$ 0.74$\times$ 0.74$\times$
RigL-DM, $S=0.10$ 46.81 $\pm$ 1.98 0.90$\times$ 0.90$\times$ 0.89$\times$
MagRan-DM, $S=0.10$ 46.00 $\pm$1.71 0.90$\times$ 0.89$\times$ 0.89$\times$
@@
Similar to Latent Diffusion, we repeated the DST experiments using the more conservative prune and regrowth rate of 0.05. The biggest improvement was seen in the Quickdraw dataset, where DST methods obtained considerably higher performances, as compared with Figure 3. Akin to the Latent Diffusion results, in higher sparsity regimes DST models mostly obtain better performance when using $p=0.05$. Please refer to Appendix F for a more in-depth analysis.
@@ 4.3 Impact of Diffusion Steps
In general, the number of sampling steps does not affect when comparing sparse and dense versions within the same number of timesteps. More experiments are presented in Appendix E. In KanjiVG, no sparse model is able to match any version of the dense model, and varying the number of timesteps appears to have little influence on the quality of the output. In CelebA-HQ, when comparing different numbers of timesteps, we observe that MagRan-DM and Static-DM with both 100 and 150 timesteps are able to outperform the Dense model using 50 timesteps. As an example, in Static-DM, $S=0.75$ with 100 timesteps vs. the dense model with 50 timesteps, Static-DM offers a theoretical speedup of 0.29$\times$ over the dense model’s Training FLOPs, and 0.57$\times$ of the Testing FLOPs, while creating better quality samples.
@@ 4.4 Limitations and Future Work
Apart from the previously mentioned computational limitations when training on CelebA-HQ and LSUN-Bedrooms, our findings demonstrate systematic trends that prompt for further investigation. Training for longer epochs could provide deeper insights into the capabilities of sparse models. Additionally, there is potential in exploring other pruning strategies and other DST hyperparameters such as $\Delta T_{e}$.
@@
LD CelebA-HQ 32.74 RigL-DM (25%, 0.5) 32.12 0.67$\times$
LD Bedrooms 31.09 MagRan-DM (10%, 0.05) 25.12 0.97$\times$
LD Imagenette 123.42 MagRan-DM (25%, 0.5) 117.32 0.91$\times$
CD QuickDraw 29.78 RigL-DM (25%, 0.05) 23.91 0.75$\times$
CD Kanji-VG 21.10 MagRan-DM (50%, 0.05) 20.32 0.51$\times$
CD VMNIST 44.21 MagRan-DM (10%, 0.5) 46.00 0.89$\times$
@@ 5 Conclusion
Open Science: Our code and trained models will be available upon publication.
Impact Statement
Our research into the efficiency of DMs holds promise towards a positive social impact, as it can allow a reduction in the required amount of computational resources, leading to lower energy consumption and a smaller carbon footprint. In addition, increased efficiency can democratize access to better generative models, making them more accessible to not only researchers, but also the general public. While this can foster fast innovation and progress, it also poses some risks, as generative models can be exploited for nefarious purposes such as spreading misinformation or manipulating public opinion.
@@ Appendix A Hardware and Software Support
There is a growing trend towards developing hardware that better supports sparse operations. In 2021, NVIDIA released the A100 GPU, which supports accelerating operations in a 2:4 sparsity pattern. Several works have already leveraged this feature [Zhou et al., 2021, Wang et al., 2024]. In order to use this capability, the sparse matrices must follow a specific structure: among each group of four contiguous values, two values must be zero, thereby fixing the sparsity level at 50%. While this structure enables significant acceleration, it supports only one static sparsity level, and makes it impossible to vary the sparsity ratio between layers.
More recently, Cerebras introduced the CS-3 AI accelerator [Lie, 2022], capable of accelerating sparse training and supporting unstructured sparsity. Using Cerebras’ CS-3 AI to accelerate training, and Neural Magic’s inference server to accelerate inference, Agarwalla et al. [2024] trained an accurate sparse Llama-2 7B model. Its accelerated training closely matched the theoretical speedup, while achieving 91.8% accuracy recovery of Llama Evaluation metrics, with 70% sparsity. This significant finding underscores the potential of sparse training to produce more efficient neural networks in practice, not just in theory.
In parallel, there have also been advancements in creating software implementations that support truly sparse-to-sparse neural network training, mostly for supervised learning tasks [Liu et al., 2021b, Curci et al., 2022]. In addition, a sparse-to-sparse denoising autoencoder has been developed by Atashgahi et al. [2022], to perform fast and robust feature selection.
@@ B.1 Model Architectures
In Latent Diffusion experiments, the model architecture is the same for LSUN-Bedrooms, CelebA-HQ and Imagenette datasets. The DM follows the architecture proposed by Rombach et al. [2021]. For the autoconder, we utilize a pre-trained model released by the Latent Diffusion authors on the project’s GitHub, [^2] https://github.com/CompVis/latent-diffusion with spatial size 64x64x3, VQ-reg regularization, and downsampling factor $f=4$.
In ChiroDiff experiments, we adopt the architecture proposed by Das et al. [2023]. The backbone network is a bidirectional GRU encoder with 3 layers, with 96 hidden units for KanjiVG, and 128 hidden units for QuickDraw. For VMNIST, the backbone network is a 2-layer bidirectional GRU encoder with 48 hidden units. We also use the code available on the project’s GitHub repository. [^3] https://github.com/dasayan05/chirodiff
@@ B.2 Choice of Datasets
We evaluated Latent Diffusion on LSUN-Bedrooms and CelebA-HQ, following their use in the original paper. Additionally, we included Imagenette, a subset of the popular ImageNet [Deng et al., 2009] dataset. For ChiroDiff, we used the same datasets evaluated as the original study: QuickDraw, KanjiVG and VMNIST. While the authors of ChiroDiff analysed seven categories of QuickDraw, namely {cat, crab, mosquito, bus, fish, yoga, flower}, we opted to reduce the number of categories to {cat, crab, mosquito} given the large number of experiments involved in our investigation. In Appendix C below we demonstrate that dataset size does not change the main outcomes. Ultimately, our goal is to compare and contrast sparse and dense models, independent of dataset size.
@@ B.4 FID calculation
For Latent Diffusion, FID is calculated using the torch-fidelity Python package, and estimated based on 10k samples and the entire training set, as in the original work. For ChiroDiff, following the original paper, we plot and save the chirographic sequences as images, and calculate the FID using the inception model provided by Ge et al. [2020], pre-trained on the QuickDraw dataset, using 10k generated samples and 20k real samples.
@@ Appendix C Experiments using the Full CelebA-HQ Dataset
Table 4: Comparison of FID scores for Latent Diffusion on CelebA-HQ using full dataset vs. reduced dataset. Results are based on the first run. Sparse models that outperform their Dense version are marked in bold. The top-performing sparse model is underlined.
Methods FID ($\downarrow$)
Full dataset Reduced dataset
Dense 32.20 29.68
Static-DM, $S=0.50$ 29.71 29.91
RigL-DM, $S=0.50$ 30.98 30.82
MagRan-DM, $S=0.50$ 26.70 30.71
@@ Appendix D Experiments using the Full ImageNet-1k Dataset
Table 5: FID score for Latent Diffusion on ImageNet.
Methods FID ($\downarrow$)
Dense 63.95
MagRan-DM, $S=0.50$ 77.39
Appendix E Experiments using Various Diffusion Timesteps

### v2

@@ 3.3.2 Experimental Details
To be able to compare the performance of different methods and different sparsity levels, we train the models for a predefined amount of epochs: 150 for Latent Diffusion datasets, and 600 for ChiroDiff datasets. For a complete description of training details please refer to Section B.3. Given the extensive number of experiments we conducted, we opted for a shorter training regime. For sampling, we use DDIM sampling (Song et al., 2021) with 100 steps for Latent Diffusion, and 50 steps for ChiroDiff, following the guidance provided in the original papers.
For completeness, we also conducted some experiments on conditional DMs, specifically on class-conditional ImageNet and class-conditional QuickDraw. See Appendix E for details on this setup.
@@ 3.3.3 Evaluation Metrics
We follow common practice and calculate the FID score (Heusel et al., 2017) to assess the performance of all models. Refer to Section B.4 for more information on FID calculation. For evaluation completeness, we also report the Kernel Inception Distance (KID) (bińkowski2018demystifying), with results presented in Appendix I. To evaluate the computational savings of the sparse methods, we report the network size (number of parameters) as a proxy for memory requirement, and the FLOPs, to estimate the computational cost of training and inference. We follow the method of FLOPs calculation described by Evci et al. (2020).
@@ 4 Experimental Results
We analyze the performance of Static-DM, MagRan-DM, and RigL-DM across various sparsity levels, and compare the results against the original Dense baseline. Additionally, we perform experiments regarding the training dynamics of Dense vs sparse models. Later on, in Section 4.4 we present experiments comparing a selection of DST vs. Dense models across various diffusion timesteps. Examples of the generated samples can be found in Appendix J. Results from the class-conditional experiments are provided in Appendix E.
@@ 4.1 Latent Diffusion
[…]
@@
[…]
@@
[…]
@@
In all datasets, Static-DM has better performance than the dynamic methods in higher sparsity setups, $S>0.5$. This is interesting, as it departs from the usual patterns found in sparse-to-sparse training for supervised learning applications and even other generative models such as GANs, where DST usually outperforms SST (Mocanu et al., 2018; Liu et al., 2023b). Liu et al. (2021c) found that, in image classification tasks, DST models consistently achieve better performance over SST with appropriate parameter exploration, i.e., exploration frequency $\Delta T_{e}$ and prune and regrowth ratio $p$. To provide insights on the importance of $p$ for DST experiments, we conducted an experiment using a prune and regrowth rate $p\in\{0.05,0.1,0.2,0.3,0.5\}$. The results are provided in Figure 8 in Appendix H. The best results were obtained with $p=0.05$.
Following this experiment, we repeated all experiments for DST methods presented in Figure 2, using $p=0.05$, and show the results in Figure 9 and Appendix H. One particularly interesting finding is that, in high sparsity regimes, such as $S=0.9$ and $S=0.75$, DST methods have consistently better performance when $p=0.05$, even outperforming Static-DM. However, this performance advantage disappears when using the more aggressive prune and regrowth rate of $p=0.5$. Please refer to Appendix H for a more in-depth analysis.
@@ 4.2 ChiroDiff
Figure 3 shows the FID scores of the studied sparse methods for ChiroDiff. For QuickDraw, we observe that both Static-DM and RigL-DM exhibit variations around the performance of the Dense model, with only a subtle tendency to deteriorate as sparsity increases. MagRan-DM consistently matches the FID of the Dense model, and is able to outperform it at 90% sparsity. These results suggest that this model is overparameterized, which would explain why it benefits significantly from sparsity, even when removing 90% of the weights. We repeated the same experiments using a smaller version of the model, the results of which are shown in Appendix F. Although most sparse models still perform comparably to the dense version, the trend of diminishing performance with increased sparsity is apparent, in contrast to the results observed with the larger model.
@@
[…]
@@
[…]
@@
[…]
@@
Similar to Latent Diffusion, we repeated the DST experiments using the more conservative prune and regrowth rate of 0.05. The biggest improvement was seen in the Quickdraw dataset, where DST methods obtained considerably higher performances, as compared with Figure 3. Akin to the Latent Diffusion results, in higher sparsity regimes DST models mostly obtain better performance when using $p=0.05$. Please refer to Appendix H for a more in-depth analysis.
4.3 Training Dynamics
In addition to final FID scores, we analysed training dynamics to better understand the behaviour of sparse models during training. In order to perform this analysis, we selected the sparsity level that achieved the overall best results in previous experiments, $S=0.25$, and plotted the FID scores across several epochs.
In general, sparse models appear to follow the trend of the corresponding dense model, which appears to indicate that they retain the stable training behaviour of dense diffusion models.
Figure 4: Comparison of training dynamics between Dense, and Static-DM, MagRan-DM and RigL-DM with $S=0.25$, for Latent Diffusion.
Figure 5: Comparison of training dynamics between Dense, and Static-DM, MagRan-DM and RigL-DM with $S=0.25$, for ChiroDiff.
@@ 4.4 Impact of Diffusion Steps
[…]
@@ 4.5 Limitations and Future Work
Apart from the previously mentioned computational limitations when training on CelebA-HQ and LSUN-Bedrooms, our findings demonstrate systematic trends that prompt for further investigation. Training for longer epochs could provide deeper insights into the capabilities of sparse models. Additionally, there is potential in exploring other pruning strategies and other DST hyperparameters such as $\Delta T_{e}$. Another interesting direction is to adjust DST hyperparameters based on the training phase, in response to changes in training dynamics. Furthermore, employing multiple sparsity masks with varying sparsity levels and dynamically changing them during training, according to the denoising timestep, is a promising line of research.
@@
[…]
@@ 5 Conclusion
Open Science: Our code and models are available at https://github.com/iclbo/sparse_to_sparse_diffusion
Acknowledgements
Research supported by Fonds National de la Recherche Luxembourg - FNR (SCRIPTOR project, grant AFR/22/17177001) and the European Innovation Council Pathfinder program (SYMBIOTIK project, grant 101071147).
The experiments presented in this paper were carried out using the HPC facilities of the University of Luxembourg (https://hpc.uni.lu) and Luxembourg’s national supercomputer MeluXina. The authors gratefully acknowledge the ULHPC and LuxProvide teams for their expert support.
@@ Appendix A Hardware and Software Support
There is a growing trend towards developing hardware that better supports sparse operations. In 2021, NVIDIA released the A100 GPU, which supports accelerating operations in a 2:4 sparsity pattern. Several works have already leveraged this feature (Zhou et al., 2021; Wang et al., 2024). In order to use this capability, the sparse matrices must follow a specific structure: among each group of four contiguous values, two values must be zero, thereby fixing the sparsity level at 50%. While this structure enables significant acceleration, it supports only one static sparsity level, and makes it impossible to vary the sparsity ratio between layers.
More recently, Cerebras introduced the CS-3 AI accelerator (Lie, 2022), capable of accelerating sparse training and supporting unstructured sparsity. Using Cerebras’ CS-3 AI to accelerate training, and Neural Magic’s inference server to accelerate inference, Agarwalla et al. (2024) trained an accurate sparse Llama-2 7B model. Its accelerated training closely matched the theoretical speedup, while achieving 91.8% accuracy recovery of Llama Evaluation metrics, with 70% sparsity. This significant finding underscores the potential of sparse training to produce more efficient neural networks in practice, not just in theory.
In parallel, there have also been advancements in creating software implementations that support truly sparse-to-sparse neural network training, mostly for supervised learning tasks (Liu et al., 2021b; Curci et al., 2022). In addition, a sparse-to-sparse denoising autoencoder has been developed by Atashgahi et al. (2022), to perform fast and robust feature selection.
@@ B.1 Model Architectures
In Latent Diffusion experiments, the model architecture is the same for LSUN-Bedrooms, CelebA-HQ and Imagenette datasets. The DM follows the architecture proposed by Rombach et al. (2021). For the autoconder, we utilize a pre-trained model released by the Latent Diffusion authors on the project’s GitHub, [^2] https://github.com/CompVis/latent-diffusion with spatial size 64x64x3, VQ-reg regularization, and downsampling factor $f=4$.
In ChiroDiff experiments, we adopt the architecture proposed by Das et al. (2023). The backbone network is a bidirectional GRU encoder with 3 layers, with 96 hidden units for KanjiVG, and 128 hidden units for QuickDraw. For VMNIST, the backbone network is a 2-layer bidirectional GRU encoder with 48 hidden units. We also use the code available on the project’s GitHub repository. [^3] https://github.com/dasayan05/chirodiff
@@ B.2 Choice of Datasets
We evaluated Latent Diffusion on LSUN-Bedrooms and CelebA-HQ, following their use in the original paper. Additionally, we included Imagenette, a subset of the popular ImageNet (Deng et al., 2009) dataset. For ChiroDiff, we used the same datasets evaluated as the original study: QuickDraw, KanjiVG and VMNIST. While the authors of ChiroDiff analysed seven categories of QuickDraw, namely {cat, crab, mosquito, bus, fish, yoga, flower}, we opted to reduce the number of categories to {cat, crab, mosquito} given the large number of experiments involved in our investigation. In Appendix C below we demonstrate that dataset size does not change the main outcomes. Ultimately, our goal is to compare and contrast sparse and dense models, independent of dataset size.
@@ B.4 FID calculation
For Latent Diffusion, FID is calculated using the torch-fidelity Python package, and estimated based on 10k samples and the entire training set, as in the original work. For ChiroDiff, following the original paper, we plot and save the chirographic sequences as images, and calculate the FID using the inception model provided by Ge et al. (2020), pre-trained on the QuickDraw dataset, using 10k generated samples and 20k real samples.
@@ Appendix C Experiments using the Full CelebA-HQ Dataset
Table 4: Comparison of FID and KID scores for Latent Diffusion on CelebA-HQ using full dataset vs. reduced dataset. Results are based on the first run. Sparse models that outperform their Dense version are marked in bold. The top-performing sparse model is underlined.
Methods FID ($\downarrow$) KID ($\downarrow$)
Full dataset Reduced dataset Full dataset Reduced dataset
Dense 32.20 29.68 0.0259 0.0241
Static-DM, $S=0.50$ 29.71 29.91 0.0242 0.0243
RigL-DM, $S=0.50$ 30.98 30.82 0.0250 0.0254
MagRan-DM, $S=0.50$ 26.70 30.71 0.0208 0.0253
@@ Appendix D Experiments using the Full ImageNet-1k Dataset
Table 5: FID and KID scores for Latent Diffusion on ImageNet.
Methods FID ($\downarrow$) KID ($\downarrow$)
Dense 63.95 0.0538
MagRan-DM, $S=0.50$ 77.39 0.0714
Appendix E Experiments using Conditional Models
While the main focus of this paper is investigating how sparse-to-sparse training affects unconditional models, we also performed some initial experiments on conditional models to explore its applicability in this setting. Specifically we conducted experiments on class-conditional ImageNet using Latent Diffusion, and class-conditional QuickDraw using ChiroDiff.
For these experiments, models were trained for 50 epochs on ImageNet and 150 on Quickdraw. For ImageNet, all classes were utilized in training and sampling was performed using $\eta$=1.0 and a classifier-free guidance scale of 3.0. For QuickDraw, seven classes were used: cat, crab, mosquito, bus, flower, yoga and fish. Examples of generated samples can be found in Figure 21 and Figure 22 in Appendix J.
On class-conditional ImageNet, the dense model outperforms the sparse version, with a moderate gap between them. On Quickdraw, models show extremely similar performance, with the best results achieved by MagRan-DM with 50% sparsity. These results suggest that sparse-to-sparse training may also be effective for conditional models.
Table 6: FID and KID scores for Latent Diffusion on class-conditional ImageNet.
Methods FID ($\downarrow$) KID ($\downarrow$)
Dense 17.33 0.0071
MagRan-DM, $S=0.50$ 20.33 0.0088
Table 7: FID and KID scores for ChiroDiff on class-conditional QuickDraw. Sparse models that outperform their Dense version are marked in bold. The top-performing sparse model is underlined.
Methods FID ($\downarrow$) KID ($\downarrow$)
Dense 31.23 0.0327
Static-DM, $S=0.50$ 31.57 0.0310
MagRan-DM, $S=0.50$ 30.20 0.0300
RigL-DM, $S=0.50$ 32.12 0.0315
Appendix F Experiments using a smaller model
In order to explore the strong results of high sparsity models ($S=0.90$) on QuickDraw using ChiroDiff, we repeated the experiments in Figure 3, using a smaller model. In this smaller model, the backbone network is a bidirectional GRU encoder consisting of 3 layers with 96 hidden units each, compared to 128 in the larger model.
In contrast with the larger model, the trend of diminishing performance as sparsity increases, present in the other datasets, can be observed. Most sparse models perform comparably to the dense, although some performance degradation is apparent in high sparsity models.
Figure 7: FID and KID scores comparisons between Dense, and Static-DM, MagRan-DM and RigL-DM with various sparsity levels, for QuickDraw dataset, with prune and regrowth rate p = 0.5.
Appendix G Experiments using Various Diffusion Timesteps